# Continual Segmentation with Disentangled Objectness Learning and Class Recognition

### 06 Mar 2024 v1

@@ 1 Introduction
Neural nets have dominated computer vision, from coarse-grained tasks like image classification to fine-grained tasks like segmentation and detection. However, the success is only sound for single-shot learning, i.e., training once to perform well with fully annotated datasets without considering the later learning process. Continual learning aims to mimic human learning by gradually obtaining knowledge in a sequential fashion [15, 17]. Particularly, continual learning shines in dense prediction tasks, such as semantic segmentation [3, 12, 5], since the annotations are laborious to obtain, and sometimes the learned samples are inaccessible, such as the nature of stream data in autonomous driving [32] or patient privacy in medical imaging [33].
@@ 2 Related Work
Continual Segmentation. Learning new knowledge by fintuning the model is a native way, but resulting catastrophic forgetting of old knowledge [16, 29, 37]. Continual learning aims to alleviate the forgetting issue by balancing the learning of new knowledge (plasticity) and the preserving of the old one (stability). While the research focus of continual learning is on classification [22, 25, 36, 1, 11, 6, 47, 39], continual segmentation started to thrive recently [3, 12, 30, 7, 34, 53, 55, 5, 38, 48, 60, 56]. As pointed out in MiB [3], the background shift in continual segmentation aggravates the forgetting issue. Cermelli et al. [3] proposed unbiased distillation and classification to mitigate the background shift. PLOP [12] circumvented the background shift issue by feature distillation and pseudo-labeling. Based on PLOP, REMINDER [34] introduced class-weighted knowledge distillation to discriminate the new and old classes. With the same motivation, Incrementer [38] designed a class deconfusion strategy and achieved competitive performance via a Transformer-based segmenter. Recently, RL-Replay [60] set a strong replay-based baseline via reinforcement learning. Some works [4, 20] started to focus on more realistic scenarios by integrating weak supervision with continual learning. The methods above follow the per-pixel classification paradigm. Although CoMFormer [5] was the first to present mask classification in continual segmentation, it applied a standard framework with distillation and pseudo-labeling and did not leverage the benefit of objectness. In contrast, our CoMasTRe exploits objectness and decouples the continual segmentation into forgetting-resistant objectness learning and class recognition.
@@ 3.2 CoMasTRe Architecture
    Next, we randomly initialize learnable positional queries $\mathcal{E}_{\mathrm{pos}}$ and input them to the objectness learning stage (stage 1), to extract positional embeddings $\mathcal{E}_{\mathrm{pos}}$ through the mask decoder $f$ for predicting class-agnostic mask proposals $\mathcal{M}$ and objectness scores $\mathcal{S}$. The training involves bipartite matching with ground truth (Fig. 2 (b)).

### 14 Mar 2024 v2

@@ 1 Introduction
Neural nets have dominated computer vision, from object recognition [18] to fine-grained classification [29, 28], segmentation [8, 48], and detection [58]. However, the success is only sound for single-shot learning, i.e., training once to perform well with fully annotated datasets without considering the later learning process. Continual learning aims to mimic human learning by gradually obtaining knowledge in a sequential fashion [15, 17]. Particularly, continual learning shines in dense prediction tasks, such as semantic segmentation [3, 12, 5], since the annotations are laborious to obtain, and sometimes the learned samples are inaccessible, such as the nature of stream data in autonomous driving [34] or patient privacy in medical imaging [35].
@@ 2 Related Work
Continual Segmentation. Learning new knowledge by fintuning the model is a native way, but resulting catastrophic forgetting of old knowledge [16, 31, 39]. Continual learning aims to alleviate the forgetting issue by balancing the learning of new knowledge (plasticity) and the preserving of the old one (stability). While the research focus of continual learning is on classification [22, 25, 38, 1, 11, 6, 50, 41], continual segmentation started to thrive recently [3, 12, 32, 7, 36, 57, 59, 5, 40, 51, 65, 60]. As pointed out in MiB [3], the background shift in continual segmentation aggravates the forgetting issue. Cermelli et al. [3] proposed unbiased distillation and classification to mitigate the background shift. PLOP [12] circumvented the background shift issue by feature distillation and pseudo-labeling. Based on PLOP, REMINDER [36] introduced class-weighted knowledge distillation to discriminate the new and old classes. With the same motivation, Incrementer [40] designed a class deconfusion strategy and achieved competitive performance via a Transformer-based segmenter. Recently, RL-Replay [65] set a strong replay-based baseline via reinforcement learning. Some works [4, 20] started to focus on more realistic scenarios by integrating weak supervision [61, 56] with continual learning. The methods above follow the per-pixel paradigm. Alhough CoMFormer [5] was the first to present mask classification in continual segmentation, it applied a standard framework with distillation and pseudo-labeling and failed to leverage the benefit of objectness. In contrast, CoMasTRe exploits objectness and decouples continual segmentation into forgetting-resistant objectness learning and class recognition.
@@ 3.2 CoMasTRe Architecture
    Next, we randomly initialize learnable positional queries $\mathcal{Q}_{\mathrm{pos}}$ and input them to the objectness learning stage (stage 1), to extract positional embeddings $\mathcal{E}_{\mathrm{pos}}$ through the mask decoder $f$ for predicting class-agnostic mask proposals $\mathcal{M}$ and objectness scores $\mathcal{S}$. The training involves bipartite matching with ground truth (Fig. 2 (b)).